### v1

# Linear Complexity Computation of Code Distance and Minimum Size of Trapping Sets for LDPC Codes with Bounded Treewidth

Abstract
It is well known that, given $b\geq 0$, finding an $(a,b)$-trapping set with the minimum $a$ in a binary linear code is NP-hard. In this paper, we demonstrate that this problem can be solved with linear complexity with respect to the code length for codes with bounded treewidth. Furthermore, suppose a tree decomposition corresponding to the treewidth of the binary linear code is known. In that case, we also provide a specific algorithm to compute the minimum $a$ and the number of the corresponding $(a,b)$-trapping sets for a given $b$ with linear complexity. Simulation experiments are presented to verify the correctness of the proposed algorithm.
Index Terms:
low-density parity-check codes, tree decomposition, trapping sets, NP-complete.
I Introduction
Low-density parity-check (LDPC) codes [1] have been extensively studied by coding theorists and practitioners over the past decade because of their near-capacity performance when decoded using low-complexity iterative decoders. However, the error floor phenomenon at high signal-to-noise ratios (SNRs) causes the bit error rate (BER) curve to flatten, limiting further performance improvement. To address this issue, researchers have focused on identifying the underlying causes of the error floor. Extensive experimental studies [2, 3, 4, 5, 6, 7] have shown that structures known as $(a,b)$-trapping sets are the primary contributors to the error floor, where $a$ represents the number of variable nodes in the structure, and $b$ denotes the number of check nodes with an odd degree. In particular, trapping sets with small values of $a$ and $b$ are the most significant contributors to the error floor.
Identifying the smallest trapping set in an LDPC code is a crucial problem, which can be formulated as follows:
Problem: MINIMUM TRAPPING SET
Instance: A Tanner graph $G$, and an intege $b\geq 0$.
Question: What is the minimum value of $a$ for which an $(a,b)$-trapping set exists in graph $G$? And find an $(a,b)$-trapping set with minimum $a$.
In particular, when $b=0$, finding the minimum size of an $(a,0)$-trapping set is equivalent to computing the minimum distance of the LDPC code. Let $H$ be the parity-check matrix corresponding to graph $G$. If there exists an $(a,0)$-trapping set $S$ in $G$, let $\bm{x}$ be the vector where the bits corresponding to the variable nodes in $S$ are set to 1, while all other bits are set to 0. It is straightforward to verify that $H\bm{x}=\bm{0}$, and the weight of $\bm{x}$ is $a$. On the other hand, if $\bm{x}$ is a weight-$a$ vector satisfying $H\bm{x}=\bm{0}$, let $S$ be the set of variable nodes corresponding to the nonzero entries of $\bm{x}$. In the induced subgraph of $S$, each check node is connected to an even number of nodes in $S$. Therefore, $S$ is an $(a,0)$-trapping set.
The study of the MINIMUM TRAPPING SET and its special cases has a long history. As early as 1971, Dominic Welsh [8] called for the development of efficient algorithms for computing the minimum distance of a code. For nearly two decades, this problem remained unsolved and was widely recognized as an open problem[9, 10, 11, 12]. It was not until 1997 that Alexander Vardy formally proved [11] that computing the minimum distance of a binary linear code is NP-hard. The study of more generalized trapping sets originated from Richardson’s semi-analytical analysis in [2], where he demonstrated that trapping sets are the primary cause of the error floor and proposed a method for rapidly estimating the error floor based on trapping sets. Since then, researchers have recognized the importance of identifying trapping sets in LDPC codes and have made considerable efforts to develop efficient methods for detecting dominant trapping sets [13, 14, 15, 16]. Unfortunately, as proven by McGregor et al. [17, 18, 19], not only is the problem of finding a minimum-size trapping set NP-hard, but even approximating its size remains NP-hard, regardless of the sparsity of the underlying graph. This implies that, for binary linear codes, no polynomial-time algorithms exist to solve the MINIMUM TRAPPING SET problem in the worst case.
Although the MINIMUM TRAPPING SET problem has been proven to be NP-hard, it continues to attract significant attention. Researchers have been persistently working on designing efficient algorithms to reduce the coefficient of exponential complexity, enabling the identification of the minimum size of trapping sets for codes of practical length (1000 to 10000) [20, 21, 22, 23]. Several reasons underscore the importance of this problem. Firstly, based on trapping sets, the error floor of LDPC codes can be quickly estimated [2, 3], which is particularly valuable for LDPC codes with error floors too low to be accurately observed through conventional computer simulations. Secondly, by modifying the decoder to account for the influence of specific trapping set structures, the decoder’s ability to mitigate error-prone patterns can be enhanced, thereby improving the overall error-correcting performance of the code(see, e.g., [24] and [25]). Furthermore, a better understanding of trapping sets can guide the design of LDPC codes by eliminating small trapping sets to achieve a lower error floor (see, e.g., [26]).
[…]
Problem: $w$-Dominating-Set
Instance: Graph $G=(V,E)$, $k,r\in\mathbb{N}$.
Parameter: treewidth $k$
Question: Let a $w$-dominating set be a vertex subset $S$ such that for all $v\in V$, either $v\in S$ or it has at least $w$ neighbors in $S$. Decide whether $G$ has a $w$-dominating set of cardinality at most $r$.
The MINIMUM TRAPPING SET problem, as an NP-hard problem, is also expected to be solvable in polynomial time with bounded treewidth.
Our primary contribution is proving that the MINIMUM TRAPPING SET problem can be solved with linear complexity in the code length for LDPC codes with bounded treewidth $k$. Furthermore, if a tree decomposition of width $k$ for the LDPC code is known, we also provide a specific algorithm to compute the size and the number of the smallest trapping sets for a given $b\geq 0$ with linear complexity.
The paper is organized as follows. Section II introduces the fundamental concepts of trapping sets in LDPC codes and treewidth. Section III demonstrates that the minimum distance can be computed with linear complexity when the treewidth is bounded. Section IV shows that the minimum size of an $(a,b)$-trapping set and the corresponding number of such sets for a given $b$ can also be computed with linear complexity under bounded treewidth. Section V provides verification using spatially coupled LDPC codes as an example. Concluding remarks are provided in Section VI.
II PRELIMINARIES
In this section, we define several important concepts, such as trapping sets and treewidth, as well as some set operations that are essential for Sections III and IV.
II-A Trapping Set
Let $G=(\mathcal{L}\cup\mathcal{R},E)$ be a bipartite graph, or Tanner graph, corresponding to the LDPC code, where $\mathcal{L}$ is the set of variable nodes, $\mathcal{R}$ is the set of check nodes, and $E$ represents the edges between $\mathcal{L}$ and $\mathcal{R}$. For a node $v\in\mathcal{L}\cup\mathcal{R}$, its degree is the number of neighbors of $v$. For subsets $L\subset\mathcal{L}$ and $R\subset\mathcal{R}$, $G[L\cup R]$ denotes the induced subgraph of $L\cup R$ in the Tanner graph. This induced subgraph has $L\cup R$ as the set of vertices, and the set of edges is given by $\{(v,c)|v\in\mathcal{L},c\in\mathcal{R},(v,c)\in E\}$. Let $\mathcal{N}_{G}(L)$ denote the set of neighbors of $L$ in $G$, we can use $G[L]$ as a shorthand for $G[L\cup\mathcal{N}_{G}(L)]$. Next, we provide the definition of a trapping set.
Definition 1.
For a Tanner graph $G=(\mathcal{L}\cup\mathcal{R},E)$ and a non-empty subset $S\subset\mathcal{L}$, let $\Gamma_{o}(G[S])$ denote the set of check nodes with odd degrees in the induced subgraph $G[S]$. We define $S$ as an $(a,b)$-trapping set in $G$ if $|S|=a$ and $|\Gamma_{o}(G[S])|=b$. The integer $a$ is referred to as the size of the trapping set $S$.
II-B Treewidth
Tree decomposition and treewidth are important concepts in graph theory. We first provide their definitions for general graphs and then apply them to Tanner graphs.
Let $G=(V,E)$ be a graph, where $V$ is the set of vertices and $E$ is the set of edges. Let $T$ be a tree. The family $(B_{t})_{t\in T}$ is a collection of vertex subsets $B_{t}\subset V$ indexed by the nodes of $T$. Then, tree decomposition can be defined as follows:
Definition 2.
The tree decomposition of $G=(V,E)$ is a pair $(T,(B_{t})_{t\in T})$ such that:
  1. 1.
    $V=\cup_{t\in T}B_{t}$;
  2. 2.
    for every edge $(u,v)\in E$, there exists a $t\in T$ such that $u\in B_{t}$ and $v\in B_{t}$;
  3. 3.
    for every vertex $v$, the set $B^{-1}(v)=\{t\in T|v\in B_{t}\}$ is non-empty and connected in $T$.
The width of a tree decomposition $(T,(B_{t})_{t\in T})$ is defined as
$\max_{t\in T}|B_{t}|-1.$
The treewidth of a graph $G$, denoted as $tw(G)$, is the minimum width among all tree decompositions of $G$. The path-width of a graph $G$, denoted as $pw(G)$, is the minimum width among all tree decompositions of $G$ in which $T$ is a path.
As defined in Definition 2, for each graph $G=(V,E)$, there exists a tree decomposition $(T,(B_{t})_{t\in T})$ where $T$ contains only a single node $t$ such that $B_{t}=V$. This implies that the treewidth of $G$ is at most |V|1|V|-1. A rooted tree decomposition $(T,(B_{t})_{t\in T})$ is a tree decomposition where $T$ has a single root node. For example, Fig.1 illustrates two rooted tree decompositions on a bipartite graph.
To better analyze algorithms based on tree decomposition, the concept of rooted nice tree decomposition [34] is essential.
Definition 3.
We say that the rooted tree decomposition $(T,(B_{t})_{t\in T})$ is nice if it satisfies the following properties:
  1. 1.
    every node in $T$ has at most two child nodes;
  2. 2.
    for the root node or leaf node $t\in T$, $B_{t}=\emptyset$;
  3. 3.
    if a node $t\in T$ has two child nodes $t_{1}$ and $t_{2}$, then $B_{t}=B_{t_{1}}=B_{t_{2}}$. node $t$ is called a join node;
  4. 4.
    if a node $t\in T$ has a single child node $t_{1}$, then one of the following two conditions must hold:
    • There exists a vertex $v\in V$ such that $B_{t}=B_{t_{1}}\cup\{v\}$. In this case, node $t$ is called an introduced node;
    • There exists a vertex $v\in V$ such that $B_{t}=B_{t_{1}}\setminus\{v\}$. In this case, node $t$ is called a forget node.
It is not hard to transform a given tree decomposition into a rooted nice tree decomposition(see [31, Lemma 13.1.3]). As shown in Fig.2, it is a rooted nice tree decomposition transformed from Fig.1(b), with the same width preserved. In the following discussion of tree decompositions in this paper, we assume that the tree decomposition is in rooted nice form.
The above definition and discussion also apply to the case where $G$ is a bipartite graph. In particular, we have the definition as follows:
Definition 4.
Let $G=(\mathcal{L}\cup\mathcal{R},E)$ be a Tanner graph with a tree decomposition $(T,(B_{t})_{t\in T})$.
  1. 1.
    For each bag $B_{t}$ with $t\in T$, we define $B_{t}^{c}$ to denote the set of check nodes in $B_{t}$, and $B_{t}^{v}$ to denote the set of variable nodes in $B_{t}$.
  2. 2.
    For each node $t\in T$, we define $T_{t}$ as the induced subgraph of $T$ consisting of all descendants of $t$ in $T$, and $G_{t}=G[\cup_{i\in T_{t}}B_{i}]$.
  3. 3.
    For each node $t\in T$, we define $\mathcal{L}(G_{t})$ and $\mathcal{R}(G_{t})$ as the sets of variable nodes and check nodes in $G_{t}$, respectively.
From the above definitions, we have $\mathcal{L}(G_{t})=\cup_{i\in T_{t}}B_{i}^{v}$ and $\mathcal{R}(G_{t})=\cup_{i\in T_{t}}B_{i}^{c}$.
(a) A bipartite graph $G$ and a tree decomposition of $G$ with width 3.
(b) A tree decomposition of $G$ with width 2.
Figure 1: We depict a bipartite graph $G$ with $tw(G)=2$, along with its two tree decompositions. Different tree decompositions may have different widths, but we are primarily concerned with the tree decomposition of minimum width.
Figure 2: The rooted nice tree decomposition corresponding to the bipartite graph in Fig.1(a). This nice tree decomposition can be obtained by adding some “auxiliary nodes” (dashed lines) to the tree decomposition in Fig.1(b).
II-C Set Operations
Let $A$ and $B$ be two sets. The fundamental set operations, including power set, intersection, union, difference, and symmetric difference, are defined as follows:
  • Power Set $(2^{A})$: The power set of $A$, denoted as $2^{A}$, is the set of all subsets of $A$, including the empty set and $A$ itself:
    $2^{A}=\{X\mid X\subseteq A\}.$
  • Intersection ($\cap$): The intersection of $A$ and $B$, denoted as $A\cap B$, is the set of elements that belong to both $A$ and $B$, i.e.,
    $A\cap B=\{x\mid x\in A\text{ and }x\in B\}.$
  • Union ($\cup$): The union of $A$ and $B$, denoted as $A\cup B$, is the set of elements that belong to at least one of $A$ or $B$, i.e.,
    $A\cup B=\{x\mid x\in A\text{ or }x\in B\}.$
  • Difference ($\setminus$): The difference of $A$ and $B$, denoted as $A\setminus B$, is the set of elements that belong to $A$ but not to $B$, i.e.,
    $A\setminus B=\{x\mid x\in A\text{ and }x\notin B\}.$
  • Symmetric Difference ($\oplus$): The symmetric difference of $A$ and $B$, denoted as $A\oplus B$, is the set of elements that belong to either $A$ or $B$, but not both, i.e.,
    $\begin{split}A\oplus B&=(A\setminus B)\cup(B\setminus A)\\ &=\{x\mid x\in A\text{ or }x\in B,\text{ but }x\notin A\cap B\}.\end{split}$
Let $U$ and $W$ be two sets. These operations satisfy several fundamental properties:
Commutativity:
$A\cup B=B\cup A,\quad A\cap B=B\cap A,\quad A\oplus B=B\oplus A.$
associativity:
$(A\cup B)\cup U=A\cup(B\cup U),\quad(A\cap B)\cap U=A\cap(B\cap U).$
Distributive Laws:
$A\cap(B\cup U)=(A\cap B)\cup(A\cap U),$
$A\cup(B\cap U)=(A\cup B)\cap(A\cup U).$
Other operations involving $\oplus$ and $\setminus$ are more complex. Here, we list several operational rules that will be used in subsequent sections:
  • $(A\oplus B)\cap U=(A\cap U)\oplus(B\cap U)$,
  • $(A\setminus B)\cap U=(A\cap U)\setminus(B\cap U)$,
  • $(A\setminus B)\cap(U\setminus W)=(A\cap U)\setminus(B\cap W)$.
Figure 3: A Venn diagram illustrating the symmetric difference and intersection of sets $R_{2}$ and $R_{3}$.
III Computer the Size and Number of the Smallest Trapping Set with $b=0$ in Codes with Bounded Treewidth
Consider a bipartite graph $G=(\mathcal{L}\cup\mathcal{R},E)$ corresponding to the LDPC code $\mathcal{C}$, with treewidth $k$. Next, for a given $b$, we will prove that finding an $(a,b)$-trapping set with the minimum $a$ in $G$ can be solved in linear complexity with respect to the code length. In this section, we focus on the case where $b=0$, while the generalization to $b>0$ will be presented in Section IV. The motivation for first considering the case $b=0$ is twofold. On the one hand, the $(a,0)$-trapping set is the most influential structure affecting the error floor of LDPC codes, and finding the smallest trapping set with $b=0$ in $G$ has received significant attention. On the other hand, the case $b>0$ is more complex. To enhance clarity and facilitate understanding, we start with the simpler case of $b=0$. Furthermore, we also show that the number of smallest trapping sets with $b=0$ contained in $G$ can be computed in linear complexity with respect to the code length.
III-A Several Important Properties of Tree Decomposition
To better utilize tree decomposition in designing algorithms for finding the smallest trapping set, we introduce several important properties to deepen our understanding of tree decomposition.
Consider a rooted nice tree decomposition of $G$, denoted by $(T,(B_{t})_{t\in T})$. Let $t$ be a node in $T$. The concepts of $G_{t}$, $B_{t}^{c}$, $B_{t}^{v}$, and other related concepts are provided in Definition 4.
Property 1: Let $TS,TS_{1},TS_{2}$ be three trapping sets in $G_{t}$, where $TS=TS_{1}\cup TS_{2}$ and $TS_{1}\cap TS_{2}=\emptyset$. Then,
$\Gamma_{o}(G_{t}[TS])=\Gamma_{o}(G_{t}[TS_{1}])\oplus\Gamma_{o}(G_{t}[TS_{2}]),$
$\Gamma_{o}(G_{t}[TS_{1}])=\Gamma_{o}(G_{t}[TS])\oplus\Gamma_{o}(G_{t}[TS_{2}]),$
$\Gamma_{o}(G_{t}[TS_{2}])=\Gamma_{o}(G_{t}[TS])\oplus\Gamma_{o}(G_{t}[TS_{1}]).$
Property 2: Let $S$ be a trapping set in $G_{t}$, and let $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are the subset of $\mathcal{R}(G_{t})$ such that $\mathcal{R}_{1}\cup\mathcal{R}_{2}=\mathcal{R}(G_{t})$, and $\mathcal{R}_{1}\cap\mathcal{R}_{2}=\emptyset$. Then,
$\begin{split}\Gamma_{o}(G_{t}[S])&=\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{1}])\cup\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{2}])\\ &=\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{2}]).\end{split}$
Property 3: Let $L\subset\mathcal{L}(G_{t})$, $R\subset\mathcal{R}(G_{t})$ be two subsets. Then $G_{t}[L\cup R]=G[L\cup R].$
Property 4: Let $t\in T$ be a join node with two child nodes $t^{\prime}$, where $\cup_{i\in T_{t}}B_{i}^{c}=\cup_{i\in T_{t^{\prime}}}B_{i}^{c}$. $S$ is a trapping set in both $G_{t}$ and $G_{t^{\prime}}$. Then,
$\Gamma_{o}(G_{t}[S])=\Gamma_{o}(G_{t^{\prime}}[S]).$
Let $t\in T$ be a join node with two child nodes $t_{1}$ and $t_{2}$, where $B_{t}=B_{t_{1}}=B_{t_{2}}$. Then,
Property 5: $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\cap\left(\cup_{i\in T_{t_{2}}}B_{i}\right)=B_{t}.$
Property 6: for any nodes $u\in\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$ and $v\in\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}$, there cannot exist an edge between $u$ and $v$ in $G$.
Property 7: $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$, $B_{t}$, and $\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}$ are three mutually disjoint sets whose union equals $\cup_{i\in T_{t}}B_{i}$.
The proofs of the first two properties mainly rely on the definition of trapping sets and set operations, while the proof of Property 3 primarily depends on the definition of $G_{t}$. The last three properties can be directly derived from the definition of tree decomposition, and thus, this proof is omitted here. Property 4 can be derived from Property 3 as follows:
$\begin{split}&\Gamma_{o}(G_{t^{\prime}}[S])=\Gamma_{o}(G_{t^{\prime}}[S\cup(\cup_{i\in T_{t^{\prime}}}B_{i}^{c})])\\ &\overset{Prop.3}{=}\Gamma_{o}(G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c})])=\Gamma_{o}(G_{t}[S]).\end{split}$
III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
Before presenting the algorithms for finding the smallest trapping set, we first introduce the following lemma, which plays a critical role in the proof.
Definition 5.
Let $(T,(B_{t})_{t\in T})$ be a rooted nice tree decomposition of $G$. Consider a node $t\in T$, node sets $R\subset B_{t}^{c}$, $Q\subset B_{t}^{v}$, and a trapping set $S$ in $G$. We define the parameters of $S$ in $G_{t}$ as $(R,Q)$ in this Section if $S$ satisfies the following properties:
  1. 1.
    $S\subset\mathcal{L}(G_{t})$;
  2. 2.
    $S\cap B_{t}^{v}=Q$;
  3. 3.
    the set of check nodes with an odd degree in $G_{t}[S]$ is exactly $R$.
Lemma 1.
Let $(T,(B_{t})_{t\in T})$ be a rooted nice tree decomposition of $G$, and $t\in T$ be a join node with two child nodes $t_{1}$ and $t_{2}$, where $B_{t}=B_{t_{1}}=B_{t_{2}}$. Suppose $S_{t_{1}}$ and $S_{t_{2}}$ are trapping sets in $G_{t_{1}}$ and $G_{t_{2}}$, respectively. The parameters of $S_{t_{1}}$ in $G_{t_{1}}$ are $(R_{t_{1}},Q)$, while the parameters of $S_{t_{2}}$ in $G_{t_{2}}$ are $(R_{t_{2}},Q)$. If the trapping set $S$ is defined as $S=S_{t_{1}}\cup S_{t_{2}}$, then its corresponding parameters in $G_{t}$ are $(R,Q)$, where
$R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}]).$
Proof.
First, we compute the second parameter associated with $S$ in $G_{t}$. According to Definition 5, we have $Q=S_{t_{1}}\cap B_{t_{1}}^{v}$, and $Q=S_{t_{2}}\cap B_{t_{2}}^{v}$, thus,
$\begin{split}S\cap B_{t}^{v}&=(S_{t_{1}}\cup S_{t_{2}})\cap B_{t}^{v}\\ &=(S_{t_{1}}\cap B_{t}^{v})\cup(S_{t_{2}}\cap B_{t}^{v})\\ &=(S_{t_{1}}\cap B_{t_{1}}^{v})\cup(S_{t_{2}}\cap B_{t_{2}}^{v})\\ &=Q.\end{split}$
Next, we compute the first parameter associated with $S$ in $G_{t}$. The trapping set $S$ can be regarded as the union of three disjoint sets of variable nodes:
(1) $Q$,
(2) $S_{t_{1}}\setminus Q$, denoted as $S_{1}$,
(3) $S_{t_{2}}\setminus Q$, denoted as $S_{2}$.
Sets $Q$, $S_{1}$, or $S_{2}$ may be empty. Let $\Gamma_{o}(G_{t}[\emptyset])=\emptyset$, we can still proceed with the analysis as follows.
For part (1): The set of check nodes with odd degrees in $G_{t}[Q]$ is given by $\Gamma_{o}(G_{t}[Q])$.
For part (2): $S_{1}$ is a set of variable nodes entirely contained in $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$. According to Property 6, we can conclude that $S_{1}$ and $\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}=\mathcal{R}(G_{t})\setminus\mathcal{R}(G_{t_{1}})$ do not have any edges connecting them. According to Property 2, this implies that $\Gamma_{o}(G_{t}[S_{1}])=\Gamma_{o}(G_{t}[S_{1}\cup\mathcal{R}(G_{t_{1}})])=\Gamma_{o}(G_{t_{1}}[S_{1}])$. Furthermore, since $S_{t_{1}}=S_{1}\cup Q$, then,
$\begin{split}\Gamma_{o}(G_{t}[S_{1}])&=\Gamma_{o}(G_{t_{1}}[S_{1}])\\ &\overset{Prop.1}{=}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t_{1}}[Q])\\ &=R_{t_{1}}\oplus\Gamma_{o}(G_{t_{1}}[Q]).\end{split}$
For part (3): Similar to part (2), we obtain
$\Gamma_{o}(G_{t}[S_{2}])=R_{t_{2}}\oplus\Gamma_{o}(G_{t_{2}}[Q]).$
Then, according to Property 1, for the union of these three parts, $S$, we have
$\begin{split}\Gamma_{o}(G_{t}[S])&=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t_{1}}[Q])\\ &\oplus\Gamma_{o}(G_{t_{2}}[Q])\oplus\Gamma_{o}(G_{t}[Q]),\end{split}$ (1)
and
$\begin{split}&\Gamma_{o}(G_{t_{1}}[Q])\oplus\Gamma_{o}(G_{t_{2}}[Q])\\ &\overset{Prop.2}{=}\Gamma_{o}(G_{t_{1}}[Q\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t_{1}}[Q\cup B_{t_{1}}^{c}])\\ &\quad\quad\;\;\oplus\Gamma_{o}(G_{t_{2}}[Q\cup\mathcal{R}_{2}])\oplus\Gamma_{o}(G_{t_{2}}[Q\cup B_{t_{2}}^{c}])\\ &\overset{Prop.3}{=}\Gamma_{o}(G_{t}[Q\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\\ &\quad\quad\;\;\oplus\Gamma_{o}(G_{t}[Q\cup\mathcal{R}_{2}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\\ &\overset{(a)}{=}\Gamma_{o}(G_{t}[Q])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}]),\end{split}$ (2)
where $\mathcal{R}_{1}=(\cup_{t\in T_{t_{1}}}B_{i}^{c})\setminus B_{t_{1}}^{c}$ and $\mathcal{R}_{2}=(\cup_{t\in T_{t_{2}}}B_{i}^{c})\setminus B_{t_{2}}^{c}$. Equation (a) holds because, according to Property 7, $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, and $B_{t}^{c}$ are three mutually disjoint sets whose union forms $\cup_{i\in T_{t}}B_{i}^{c}=\mathcal{R}(G_{t})$. This result follows from Property 2.
Combining Equations (1) and (2) completes the proof of the Lemma.
Next, we present an algorithm to compute the size and the number of the smallest trapping set with $b=0$ in $G$, along with its computational complexity. Additionally, we provide a detailed proof to verify the correctness of the algorithm.
Theorem 2.
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest trapping sets with $b=0$ in $G$ can be determined in time $O(n2^{2k})$.
Proof.
A nice tree decomposition of $G$ is given by $(T,(B_{t})_{t\in T})$, with root $r$ and width $k$. For each bag $B_{t}$, we define two functions,
$f_{t}:(L\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}})\mapsto f_{t}(L,Q)\in\mathbb{Z},$
and
$g_{t}:(L\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}})\mapsto g_{t}(L,Q)\in\mathbb{Z}.$
where $f_{t}(R,Q)$ denotes the size of the smallest trapping set in $G_{t}$ with parameters $(R,Q)$, while $g_{t}(R,Q)$ represents the number of such trapping sets. Note that $\emptyset$ is not a trapping set. Therefore, even if $\emptyset\cap B_{t}^{v}=\emptyset$ and $\Gamma_{o}(G_{t}[\emptyset])=\emptyset$, we don’t have $f_{t}(\emptyset,\emptyset)=0$ and $f_{t}(\emptyset,\emptyset)=1$.
Since all and only the trapping sets with $b=0$ correspond to the parameters $(\emptyset,\emptyset)$ in $G_{r}$, it follows that once we obtain $f_{r}(\emptyset,\emptyset)$ and $g_{r}(\emptyset,\emptyset)$, we can determine the size and the number of the smallest trapping sets with $b=0$ in the Tanner graph $G$.
Leaf node: Suppose $t$ is a leaf node. Then,
$f_{t}(R,Q)=+\infty,$
and
$g_{t}(R,Q)=0,$
for all $R\in 2^{B_{t}^{c}}=\emptyset$ and $Q\in 2^{B_{t}^{v}}=\emptyset$.
It takes $O(1)$ time to compute all $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$ for each leaf node $t$. After computing $f_{t}$ and $g_{t}$ for the leaf nodes, we then visit the bags of the tree decomposition from the leaves to the root, calculating the corresponding $f$-functions and $g$-functions at each step according to the following rules.
Let $t$ be a non-leaf node and assume we have count $f_{t^{\prime}}(R,Q)$ and $g_{t^{\prime}}(R,Q)$ , for all $R\in 2^{B_{t^{\prime}}^{c}}$ and $Q\in 2^{B_{t^{\prime}}^{v}}$, where $t^{\prime}$ is a child node of $t$. We then design an algorithm for five types of $t$ to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$, for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$. We only provide a justification for the computation of $f_{t}(R,Q)$, as the validity of $g_{t}(R,Q)$ follows directly from the proof of $f_{t}(R,Q)$.
introduced variable node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{v\}$, where $v$ is a variable node. This means the functions $f_{t}$ and $g_{t}$ are extensions of the domains of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively, as $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}\cup\{v\}})$. We can compute $f_{t}$ and $g_{t}$ using the equations (3) and (4) for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$.
$f_{t}(R,Q)=\begin{cases}f_{t^{\prime}}(R,Q),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\}\text{and }R=\Gamma_{o}(G_{t}[v]);\\ f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\})+1,&\text{otherwise.}\end{cases}$ (3)
$g_{t}(R,Q)=\begin{cases}g_{t^{\prime}}(R,Q),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\}\text{and }R=\Gamma_{o}(G_{t}[v]);\\ g_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\}),&\text{otherwise.}\end{cases}$ (4)
First, we prove that the left-hand side of (3) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$.
If $(R,Q)$ satisfies $v\notin Q$, then $v\notin S$ and $S\subset\mathcal{L}(G_{t^{\prime}})$ is also a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap(B_{t}^{v}\setminus\{v\})\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S]),S\cap B_{t}^{v}\Big{)}\\ &=(R,Q)\subset\left(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}\right).\end{split}$ (5)
This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
If $(R,Q)$ satisfies $Q=\{v\}$ and $R=\Gamma_{o}(G_{t}[v])$, then $S$ is either $\{v\}$ or $\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set with $b=0$ in $G_{t}^{\prime}$ such that $S_{t^{\prime}}\cap B_{t^{\prime}}=\emptyset$. In this case, we have $f_{t}(R,Q)=1$ and $g_{t}(R,Q)=1$.
For other cases of $(R,Q)$, $S_{t}$ has the form $S=\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S_{t^{\prime}}$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S_{t^{\prime}}]),S_{t^{\prime}}\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S\setminus\{v\}]),(S\setminus\{v\})\cap(B_{t}^{v}\setminus\{v\})\Big{)}\\ &\overset{Prop.1}{=}\Big{(}\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G_{t}[v]),(S\cap B_{t}^{v})\setminus(\{v\}\cap\{v\})\Big{)}\\ &\overset{(b)}{=}\Big{(}R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\}\Big{)}\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}).\end{split}$ (6)
The reason why $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$ holds in (b) is that $v$ belongs only to $B_{t}$. According to the definition of tree decomposition, $v$ can only be adjacent to vertices in $B_{t}^{c}$ within $G_{t}$. This leads to $\mathcal{N}_{G_{t}}(v)\subset B_{t}^{c}$, which results $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$. Therefore, $|S|=|S_{t^{\prime}}|+1$ implies $f_{t}(R,Q)\geq f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\})+1$.
On the other hand, we now proceed to prove that the left-hand side of (3) is less than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, if $(R,Q)$ satisfies $v\notin Q$, let $R^{\prime}=R\subset B_{t^{\prime}}^{c}$ and $Q^{\prime}=Q\subset B_{t^{\prime}}^{v}$, and $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$. Then, $S_{t^{\prime}}$ remains a trapping set in $G_{t}$. By reversing the steps in (5), we obtain that $S_{t^{\prime}}$ retains the parameters $(R^{\prime},Q^{\prime})$ in $G_{t}$, which leads to $f_{t}(R^{\prime},Q^{\prime})\leq f_{t^{\prime}}(R^{\prime},Q^{\prime})$. If $(R,Q)$ satisfies $Q=\{v\}$ and $R\neq\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$ or $v\in Q\neq\{v\}$, let $R^{\prime}=R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])\subset B_{t^{\prime}}^{c},Q^{\prime}=Q\backslash\{v\})\subset B_{t^{\prime}}^{v}$, $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$, and $S=\{v\}\cup S_{t^{\prime}}$ be a trapping set in $G_{t}$. By applying the reverse derivation of (6), we obtain that the parameters of $S$ in $G_{t}$ remain as $(R^{\prime}\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q^{\prime}\cup\{v\})$. This leads to $f_{t^{\prime}}(R^{\prime},Q^{\prime})+1\geq f_{t}(R^{\prime}\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q^{\prime}\cup\{v\}).$ Thus, we have $f_{t}(R,Q)\leq f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\})+1.$
It takes $O(2^{k})$ time for each introduced variable node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Forget variable node: Suppose $t$ is a forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{v\}$, where $v$ is a variable node. This means the functions $f_{t}$ and $g_{t}$ are restrictions of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively. Specifically, $f_{t}$ and $g_{t}$ are defined over a smaller domain $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}\setminus\{v\}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\},$ (7)
and
$\begin{split}g_{t}(R,Q)&=\mathbb{1}_{f_{t}(R,Q)=f_{t^{\prime}}(R,Q)}g_{t}(R,Q)\\ &+\mathbb{1}_{f_{t}(R,Q)=f_{t^{\prime}}(R,Q\cup\{v\})}g_{t}(R,Q\cup\{v\}),\end{split}$ (8)
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where $\mathbb{1}_{x=y}$ is the indicator function,
$\mathbb{1}_{x=y}=\begin{cases}1,&\text{if $x=y$};\\ 0,&\text{if $x\neq$y}.\end{cases}$ (9)
Next, we provide proof for the validity of Equation (7). For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S],S\cap B_{t^{\prime}}^{v}\Big{)}\\ &=\Big{(}G_{t^{\prime}}[S],S\cap(B_{t}^{v}\cup\{v\})\Big{)}\\ &\overset{Prop.4}{=}\Big{(}G_{t}[S],(S\cap B_{t}^{v})\cup(S\cap\{v\})\Big{)}\\ &=\Big{(}R,Q\cup(S\cap\{v\})\Big{)}.\end{split}$ (10)
If $v\notin S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. If $v\in S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q\cup\{v\})\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. Therefore,
$f_{t}(R,Q)\geq\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\}.$
Conversely, for any trapping set $S_{t^{\prime}}$ in $G_{t^{\prime}}$ with parameters $(R,Q)$ or $(R,Q\cup\{v\})$, its parameters in $G_{t}$ can be derived as $(R,Q)$ by applying Equation (10). This implies that
$f_{t}(R,Q)\leq\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\}.$
It takes $O(2^{k})$ time for each forget variable node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
introduced check node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{c\}$, where $c$ is a check node. This means the functions $f_{t}$ and $g_{t}$ are extensions of the domains of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively, as $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}\cup\{c\}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\begin{cases}f_{t^{\prime}}(R\setminus\{c\},Q),&\text{if }c\in R\text{, $E(c,Q)$ is odd;}\\ f_{t^{\prime}}(R,Q),&\text{if }c\notin R\text{, $E(c,Q)$ is even;}\\ +\infty,&\text{otherwise},\end{cases}$ (11)
and
$g_{t}(R,Q)=\begin{cases}g_{t^{\prime}}(R\setminus\{c\},Q),&\text{if }c\in R\text{, $E(c,Q)$ is odd;}\\ g_{t^{\prime}}(R,Q),&\text{if }c\notin R\text{, $E(c,Q)$ is even;}\\ 0,&\text{otherwise},\end{cases}$ (12)
for all $R\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$, where $E(c,Q)$ denotes the number of edges between $c$ and $Q$ in $G_{t}$, and it is easy to compute as it only requires determining the degree of $c$ in $G_{t}[Q\cup\{c\}]$. Specifically, $E(c,Q)$ being even includes the case where $E(c,Q)=0$.
First, we prove that the left-hand side of (11) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. Since $\mathcal{L}(G_{t})=\mathcal{L}(G_{t^{\prime}})$, $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c}\setminus\{c\})]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.2}{=}\Big{(}\Gamma_{o}\big{(}G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c})]\big{)}\oplus\Gamma_{o}\big{(}G_{t}[S\cup\{c\}]\big{)},Q\Big{)}\\ &\overset{(c)}{=}\Big{(}\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G[Q\cup\{c\}]),Q\Big{)}\\ &=\Big{(}R\oplus\Gamma_{o}(G[Q\cup\{c\}]),Q\Big{)}.\end{split}$ (13)
The validity of equation (c) follows from the definition of tree decomposition, which implies that $c$ can only be adjacent to variable nodes in $B_{t}^{v}$ and not to those in $S\setminus B_{t}^{v}$. Consequently, the number of edges between $c$ and $S$ is equal to the number of edges between $c$ and $S\cap B_{t}^{v}=Q$.
If $c$ and $S$ have an odd number of edges in $G_{t}$, then $c\in R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\{c\}$. In this case, the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R\setminus\{c\},Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R\setminus\{c\},Q)$.
If $c$ and $S$ have an even number of edges in $G_{t}$, then $c\notin R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\emptyset$. Consequently, the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
There are no other possible cases of trapping sets in $G_{t}$. Hence, $f_{t}(R,Q)=\infty$ for any other cases of $R$ and $Q$.
As for the proof that the left-hand side of (11) is less than or equal to the right-hand side, consider each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$. If $(R,Q)$ satisfies $c\in R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\{c\}$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R\setminus\{c\},Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. By (13), $S_{t^{\prime}}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Therefore, $f_{t}(R,Q)\leq f_{t^{\prime}}(R\setminus\{c\},Q)$. If $(R,Q)$ satisfies $c\notin R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\emptyset$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. By (13), $S_{t^{\prime}}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Therefore, $f_{t}(R,Q)\leq f_{t^{\prime}}(R,Q)$.
It takes $O(2^{k})$ time for each introduced check node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Forget check node: Suppose $t$ is an forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{c\}$, where $c$ is a check node. This means the functions $f_{t}$ and $g_{t}$ are restrictions of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively. Specifically, $f_{t}$ and $g_{t}$ are defined over a smaller domain $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}\setminus\{c\}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=f_{t}^{\prime}(R,Q),$ (14)
and
$g_{t}(R,Q)=g_{t}^{\prime}(R,Q),$ (15)
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$.
Next, we provide proof for the validity of Equation (14). For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S]),S\cap B_{t}^{v}\Big{)}\\ &=(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}).\end{split}$
This implies that $f_{t}(R,Q)\leq f_{t^{\prime}}(R,Q)$. Conversely, it is also easy to verify that any trapping set in $G_{t^{\prime}}$ with parameters $(R,Q)$ retains the same parameters in $G_{t}$, which leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
Note that the reduction in the domain of the functions $f_{t}$ and $g_{t}$ in equations (14) and (15) results in the loss of certain function values. Specifically, for each pair $(R^{\prime}\subset B_{t^{\prime}}^{c},Q^{\prime}\subset B_{t^{\prime}}^{v})$ with $c\in R^{\prime}$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$. However, since $\Gamma_{o}(G_{t}[S_{t^{\prime}}])\not\subset B_{t}^{c}$, we are unable to compute the minimum size of $S_{t^{\prime}}$ using $f_{t}$.
Nevertheless, this does not affect our ability to determine the size of the smallest trapping set with $b=0$ in $G$. This is because, according to the definition of tree decomposition, $c$ is not adjacent to any variable node in $\mathcal{L}(G)\setminus\mathcal{L}(G_{t})$. Let $S=S^{\prime}\cup S_{t^{\prime}}$ be any trapping set in $G$ such that $S^{\prime}\in\mathcal{L}(G)\setminus\mathcal{L}(G_{t})$, this leads to $c\in\Gamma_{o}(G[S])$. Consequently, the trapping set $S$ cannot correspond to $b=0$, implying that whether or not we store the minimum size of $S_{t^{\prime}}$ does not affect the final result of the algorithm.
It takes $O(2^{k})$ time for each forget check node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Join node: Suppose $t$ is a join node, and $t_{1}$ and $t_{2}$ be its two child nodes with $B_{t}=B_{t1}=B_{t2}$. Assume we have count $f_{t_{1}}(R,Q)$, $g_{t_{1}}(R,Q)$, $f_{t_{2}}(R,Q)$, and $g_{t_{2}}(R,Q)$, for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$. Furthermore, this implies that these functions share the same domain as $f_{t}(R,Q)$ and $g_{t}(R,Q)$, which is $(2^{B_{t}^{c}},2^{B_{t}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\begin{cases}F_{t}(R,Q),\text{\qquad\qquad\qquad\qquad\qquad\;\;\;\; if }Q\neq\emptyset\\ \min\{F_{t}(R,Q),f_{t_{1}}(R,Q),f_{t_{2}}(R,Q)\},\text{if }Q=\emptyset\end{cases}$ (16)
$g_{t}(R,Q)=\begin{cases}G_{t}(R,Q),&\text{if }Q\neq\emptyset;\\ \mathbb{1}_{f_{t}(R,Q)=F_{t}(R,Q)}G_{t}(R,Q)\\ +\mathbb{1}_{f_{t}(R,Q)=f_{t_{1}}(R,Q)}g_{t_{1}}(R,Q)\\ +\mathbb{1}_{f_{t}(R,Q)=f_{t_{2}}(R,Q)}g_{t_{2}}(R,Q),&\text{if }Q=\emptyset,\end{cases}$ (17)
$\begin{split}&F_{t}(R,Q)=\min_{(R_{t_{1}},R_{t_{2}})}\{f_{t_{1}}(R_{t_{1}},Q)+f_{t_{2}}(R_{t_{2}},Q)-|Q|\},\end{split}$ (18)
$\begin{split}&G_{t}(R,Q)=\sum_{(R_{t_{1}},R_{t_{2}})}g_{t_{1}}(R_{t_{1}},Q)\times g_{t_{2}}(R_{t_{2}},Q),\end{split}$ (19)
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where (18) takes the minimum over all pairs $(R_{t_{1}},R_{t_{2}})$ satisfying $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$, (19) takes the sum over all pairs $(R_{t_{1}},R_{t_{2}})$ that attain the minimum value in (18) and $\mathbb{1}_{x=y}$ is the indicator function.
First, we prove that the left-hand side of (16) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ can be regarded as the union of three disjoint sets of variable nodes:
(1) $Q=S\cap B_{t}^{v}$;
(2) $S_{1}:=(S\setminus Q)\cap(\cup_{i\in T_{t_{1}}}B_{i})$;
(3) $S_{2}:=(S\setminus Q)\cap(\cup_{i\in T_{t_{2}}}B_{i})$.
Sets $Q$, $S_{1}$, or $S_{2}$ may be empty sets and $S$ is non-empty. Let $S_{t_{1}}=S_{1}\cup Q$ and $S_{t_{2}}=S_{2}\cup Q$. When both $S_{t_{1}}$ and $S_{t_{2}}$ are non-empty sets, let $S_{t_{1}}$ have parameters $(R_{t_{1}},Q)$ in $G_{t_{1}}$, and let $S_{t_{2}}$ have parameters $(R_{t_{2}},Q)$ in $G_{t_{2}}$. According to Lemma 1, we have $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$ and $|S|=|S_{t_{1}}|+|S_{t_{2}}|-|Q|$. This implies that $f_{t}(R,Q)\geq F_{t}(R,Q)$.
When $S_{t_{1}}=\emptyset$ and $S_{t_{2}}\neq\emptyset$, $S\subset(\cup_{i\in T_{t_{2}}}B_{i})$ is also a trapping set in $G_{t_{2}}$, which implies that the parameters of $S$ in $G_{t_{2}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t_{2}}[S]),S\cap B_{t_{2}}^{v}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t_{2}}[S_{2}]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.3}{=}\Big{(}\Gamma_{o}(G_{t}[S_{2}\cup(\cup_{i\in T_{t_{2}}}B_{i}^{c})]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.6}{=}\Big{(}\Gamma_{o}(G_{t}[S_{2}]),Q\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t}[S]),Q\Big{)}\\ &=(R,Q).\end{split}$ (20)
Therefore, $f_{t}{(R,Q)}\geq f_{t_{2}}(R,Q)$.
When $S_{t_{1}}\neq\emptyset$ and $S_{t_{2}}=\emptyset$, a similar proof shows that $f_{t}(R,Q)\geq f_{t_{1}}(R,Q)$.
By combining these three cases, we can prove that the left-hand side of Equation (16) is greater than or equal to the right-hand side.
On the other hand, we prove that the left-hand side of Equation (16) is less than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S_{t_{1}}$ and $S_{t_{2}}$ be any trapping sets in $G_{t_{1}}$ and $G_{t_{2}}$ with parameters $(R_{t_{1}},Q)$ and $(R_{t_{2}},Q)$, respectively, such that $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$. Let $S_{t}=S_{t_{1}}\cup S_{t_{2}}$. Then, according to Lemma 1, $S_{t}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Thus, $F_{t}(R,Q)\geq f_{t}(R,Q).$ Furthermore, when $Q=\emptyset$, let $S_{t_{j}}^{\prime}$ be any trapping set with parameters $(R,\emptyset)$ in $G_{t_{j}}$, where $j=1$ and 2. Then, $S_{t_{j}}^{\prime}$ is also a trapping set in $G_{t}$. According to (20), we have that the parameters of $S_{t_{j}}^{\prime}$ in $G_{t}$ are given by $(R,Q)$. Therefore, we have $f_{t_{j}}(R,\emptyset)\geq f_{t}(R,\emptyset)$.
For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, it takes $O(2^{k})$ time to obtain the minimum value in (18). Therefore, the computational complexity for each join node is $O(2^{2k})$, while for each other type of node, it is $O(2^{k})$. Given that a nice tree decomposition $(T,(B_{t})_{t\in T})$ of the Tanner graph $G$ contains $O(n)$ nodes in $T$, the overall time complexity of the algorithm is $O(n2^{2k})$. ∎
Note that during the proof process, the computation of $f_{t}$ and $g_{t}$ depends only on the functions $f_{t^{\prime}}$ and $g_{t^{\prime}}$, where $t^{\prime}$ is the child of $t$ in $T$. Therefore, when traversing the bags from the leaves to the root, the functions $f_{i}$ and $g_{i}$, where $i$ is a descendant of $t^{\prime}$, do not need to be stored.
By executing the algorithm in Theorem 2 several times, we can find an $(a,0)$-trapping set with minimum $a$ in $G$. This leads to the following corollary:
Corollary 3.
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. We can find an $(a,0)$-trapping sets with minimum $a$ in $G$ in time $O(n2^{2k})$.
Proof.
In fact, we can first execute the algorithm in Theorem 2 to determine the minimum size $a_{0}$ of an $(a,0)$-trapping set. Then, by running the algorithm $a_{0}$ times, we can identify all $(a_{0},0)$-trapping sets. ∎
Sometimes, computing the treewidth of a given LDPC code is challenging. However, if the pathwidth of the LDPC code is known, we can derive the following corollary by applying the algorithm in Theorem 2.
Corollary 4.
Let $G$ be a Tanner graph with path-width $pw(G)=k^{\prime}$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest trapping sets with $b=0$ in $G$ can be determined in time $O(n2^{k^{\prime}})$.
Proof.
Let $(T,(B_{t})_{t\in T})$ be a nice tree decomposition of $G$ with width $k^{\prime}$, where $T$ is a path. In this case, $T$ contains no join nodes and consists only of leaf nodes, introduced variable nodes, forget variable nodes, introduced check nodes, and forget check nodes. According to Theorem 2, each node requires $O(2^{k^{\prime}})$ time, leading to an overall computational complexity of $O(n2^{k^{\prime}})$ for the algorithm. ∎
IV Computer The Size and Number of the Smallest Trapping Set with $b>0$ in Codes with Bounded Treewidth
As a generalization of Section III, this section presents an algorithm to find an $(a,b)$-trapping set with the minimum $a$ in Tanner graph $G$, given $b>0$, in linear complexity with respect to the code length. Furthermore, we also demonstrate that the number of smallest trapping sets with $b>0$ contained in $G$ can be computed in linear complexity with respect to the code length.
First, we provide the generalized form of the algorithm in Theorem 2 to compute the size and the number of the smallest trapping sets with $b>0$ in $G$, along with its proof.
Definition 6.
Let $(T,(B_{t})_{t\in T})$ be a rooted nice tree decomposition of Tanner graph $G$. Consider a node $t\in T$, node sets $I\subset B_{t}^{c}$, $Q\subset B_{t}^{v}$, an integer $d$ and a trapping set $S$ in $G$. We define the parameters of $S$ in $G_{t}$ as $(I,Q,d)$ in this Section if $S$ satisfies the following properties:
  1. 1.
    $S\subset\mathcal{L}(G_{t})$;
  2. 2.
    $S\cap B_{t}^{v}=Q$;
  3. 3.
    $\Gamma_{o}(G_{t}[S])\cap B_{t}^{c}=I$,
  4. 4.
    $|\Gamma_{o}(G_{t}[S])\setminus I|=d$.
Theorem 5.
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. Then, for a given $b$, the size and the number of the smallest trapping sets with $b>0$ in $G$ can be determined in time $O(n2^{2k+2b})$.
Proof.
A nice tree decomposition of $G$ is given by $(T,(B_{t})_{t\in T})$, with root $r$ and width $k$. For each bag, we define two functions as follows:
$f_{t}:(I\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}},d\in\{0,1,...,b\})\mapsto f_{t}(I,Q,d)\in\mathbb{Z},$
and
$g_{t}:(I\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}},d\in\{0,1,...,b\})\mapsto g_{t}(I,Q,d)\in\mathbb{Z}.$
where $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$ denote the size and the number of the smallest trapping set in $G_{t}$ with parameters $(I,Q,d)$, respectively.
As Theorem 2, once we obtain $f_{r}(\emptyset,\emptyset,b)$ and $g_{r}(\emptyset,\emptyset,b)$, we determine the size and the number of the smallest trapping sets with $b>0$ in $G$.
Leaf node: Suppose $t$ is a leaf node. Then,
$f_{t}(I,Q,d)=+\infty,$
and
$g_{t}(I,Q,d)=0,$
for all $I\in 2^{B_{t}^{c}}=\emptyset$, $Q\in 2^{B_{t}^{v}}=\emptyset$, and $d\in\{0,1,...,b\}$.
It takes $O(b)$ time to compute all $f_{t}(I,Q,b)$ and $g_{t}(I,Q,b)$ for each leaf node $t$. Similar to Theorem 2, we traverse the bags from the leaf nodes to the root node and compute the corresponding $f$ and $g$ functions.
Let $t$ be a non-leaf node and assume we have count $f_{t^{\prime}}(I,Q,d)$ and $g_{t^{\prime}}{(I,Q,d)}$, for all $I\in 2^{B_{t^{\prime}}^{c}}$, $Q\in 2^{B_{t^{\prime}}^{v}}$ and $d\in\{0,1,...,b\}$, where $t^{\prime}$ is a child node of $t$. We design an algorithm to compute $f_{t}(I,Q,d)$ and $g_{t}{(I,Q,d)}$, for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$ and $d\in\{0,1,...,b\}$.
Note that the proof idea here is similar to that of Theorem 2, and we only emphasize the differences. Furthermore, when deriving the expressions of $f_{t}$ and $g_{t}$ in terms of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, we only prove that the left-hand side of the expression is greater than or equal to the right-hand side. The case where the left-hand side is less than or equal to the right-hand side can be derived from the proof process of this theorem and Theorem 2, and thus, it is omitted here.
introduced variable node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{v\}$, where $v$ is a variable node. This means the functions $f_{t}$ and $g_{t}$ are extensions of the domains of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively, as $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}\cup\{v\}},\{0,1,...,b\})$. We can compute $f_{t}$ and $g_{t}$ using the equations (21) and (22) for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$.
$f_{t}(I,Q,d)=\begin{cases}f_{t^{\prime}}(I,Q,d),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\},I=\Gamma_{o}(G_{t}[v]),\text{and }d=0;\\ f_{t^{\prime}}(I\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\},d)+1,&\text{otherwise,}\end{cases}$ (21)
$g_{t}(I,Q,d)=\begin{cases}g_{t^{\prime}}(I,Q,d),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\},I=\Gamma_{o}(G_{t}[v]),\text{ and }d=0;\\ g_{t^{\prime}}(I\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\},d),&\text{otherwise},\end{cases}$ (22)
Next, we prove that the left-hand side of (21) is greater than or equal to the right-hand side. For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, let $S$ be any trapping set in $G_{t}$ with parameters $(I,Q,d)$.
If $(I,Q,d)$ satisfies $v\notin Q$, then $v\notin S$ and $S\subset\mathcal{L}(G_{t^{\prime}})$ is also a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S])\cap B_{t^{\prime}}^{c},S\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem\;2}{=}\Big{(}\Gamma_{o}(G_{t}[S])\cap B_{t^{\prime}}^{c},S\cap B_{t}^{v},\Big{|}\Gamma_{o}(G_{t}[S])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t}[S])\cap B_{t}^{c},S\cap B_{t}^{v},\Big{|}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\Big{|}\Big{)}\\ &=(I,Q,d)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\}).\end{split}$ (23)
This leads to $f_{t}(I,Q,d)\geq f_{t^{\prime}}(I,Q,d)$.
If $(I,Q,d)$ satisfies $Q=\{v\}$, $I=\Gamma_{o}(G_{t}[v])$ and $d=0$, then $S$ is either $\{v\}$ or $\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set with $b=0$ in $G_{t}^{\prime}$ such that $S_{t^{\prime}}\cap B_{t^{\prime}}=\emptyset$. In this case, we have $f_{t}(I,Q,d)=1$ and $g_{t}(I,Q,d)=1$.
For other cases of $(I,Q,d)$, $S_{t}$ has the form $S=\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S_{t^{\prime}}$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S_{t^{\prime}}])\cap B_{t^{\prime}}^{c},S_{t^{\prime}}\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S_{t^{\prime}}])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem2}{=}\Big{(}(\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]))\cap B_{t}^{c},Q\setminus\{v\},\\ &\qquad\qquad\;\;\Big{|}(\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G_{t}[v]))\Big{|}\Big{)}\\ &=\Big{(}I\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\},\Big{|}(\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\Big{|}\Big{)}\\ &=\Big{(}I\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\},d\Big{)}\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\}).\end{split}$ (24)
Therefore, $|S|=|S_{t^{\prime}}|+1$ implies $f_{t}(I,Q)\geq f_{t^{\prime}}(I\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\})+1$.
The algorithm takes $O(b2^{k})$ time for each introduced variable node.
Forget variable node: Suppose $t$ is a forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{v\}$, where $v$ is a variable node. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(I,Q,d)=\min\{f_{t^{\prime}}(I,Q,d),f_{t^{\prime}}(I,Q\cup\{v\},d)\},$ (25)
and
$\begin{split}g_{t}(I,Q,d)&=\mathbb{1}_{f_{t}(I,Q,d)=f_{t^{\prime}}(I,Q,d)}g_{t}(I,Q,d)\\ &+\mathbb{1}_{f_{t}(I,Q,d)=f_{t^{\prime}}(I,Q\cup\{v\},d)}g_{t}(I,Q\cup\{v\},d),\end{split}$ (26)
for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$, where $\mathbb{1}_{x=y}$ is the indicator function.
Next, we prove that the left-hand side of (25) is greater than or equal to the right-hand side. For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, let $S$ be any trapping set in $G_{t}$ with parameters $(I,Q,d)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]\cap B_{t^{\prime}}^{c},S\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S]\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem\;2}{=}\Big{(}G_{t}[S]\cap B_{t}^{c},Q\cup(S\cap\{v\}),\Big{|}G_{t}[S]\setminus B_{t}^{c}\Big{|}\Big{)}\\ &=\Big{(}I,Q\cup(S\cap\{v\}),d\Big{)}.\end{split}$ (27)
If $v\notin S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(I,Q,d)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\})$. If $v\in S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(I,Q\cup\{v\},d)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\}))$. Therefore,
$f_{t}(I,Q,d)\geq\min\{f_{t^{\prime}}(I,Q,d),f_{t^{\prime}}(I,Q\cup\{v\},d)\}.$
It takes $O(b2^{k})$ time for each forget variable node.
introduced check node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{c\}$, where $c$ is a check node. We can compute $f_{t}$ and $g_{t}$ using the following equations:
[…]
and
[…]
for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$, where $E(c,Q)$ is defined in Theorem 2.
Next, we prove that the left-hand side of (28) is greater than or equal to the right-hand side. For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, let $S$ be any trapping set in $G_{t}$ with parameters $(I,Q,d)$. Since $\mathcal{L}(G_{t})=\mathcal{L}(G_{t^{\prime}})$, $S$ remains a trapping set in $G_{t^{\prime}}$.
If $c$ and $S$ have an odd number of edges in $G_{t}$, then $c\in I$, $\Gamma_{o}(G[Q\cup\{c\}])=\{c\}$ and $\{c\}\in\Gamma_{o}(G_{t}[S])$. In this case, the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S])\cap B_{t^{\prime}}^{c},S\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem\;2}{=}\Big{(}\big{(}\Gamma_{o}(G_{t}[S])\oplus\{c\}\big{)}\cap\big{(}B_{t}^{c}\setminus\{c\}\big{)},Q,\\ &\Big{|}\big{(}\Gamma_{o}(G_{t}[S])\oplus\{c\}\big{)}\setminus\big{(}B_{t}^{c}\setminus\{c\}\big{)}\Big{|}\Big{)}\\ &=\Big{(}\big{(}I\setminus\{c\}\big{)}\oplus\big{(}\{c\}\setminus\{c\}\big{)},Q,\\ &\Big{|}\Big{(}\big{(}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\big{)}\cup\{c\}\Big{)}\oplus\Big{(}\emptyset\cup\{c\}\Big{)}\Big{|}\Big{)}\\ &=\Big{(}I\setminus\{c\},Q,\Big{|}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\Big{|}\Big{)}\\ &=(I\setminus\{c\},Q,d)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\}).\end{split}$ (30)
This leads to $f_{t}(I,Q,d)\geq f_{t^{\prime}}(I\setminus\{c\},Q,d)$.
If $c$ and $S$ have an even number of edges in $G_{t}$, then $c\notin I$, $\Gamma_{o}(G[Q\cup\{c\}])=\emptyset$ and $c\notin\Gamma_{o}(G_{t}[S])$. Consequently, the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S])\cap B_{t^{\prime}}^{c},S\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem\;2}{=}\Big{(}\Gamma_{o}(G_{t}[S])\cap\big{(}B_{t}^{c}\setminus\{c\}\big{)},Q,\\ &\quad\quad\quad\quad\;\;\;\Big{|}\Gamma_{o}(G_{t}[S])\setminus\big{(}B_{t}^{c}\setminus\{c\}\big{)}\Big{|}\Big{)}\\ &=\Big{(}I\cap\emptyset,Q,\Big{|}\big{(}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\big{)}\cup\emptyset\Big{|}\Big{)}\\ &=(I,Q,d)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v},\{0,1,...,b\}).\end{split}$ (31)
This leads to $f_{t}(I,Q,d)\geq f_{t^{\prime}}(I,Q,d)$.
There are no other possible cases of trapping sets in $G_{t}$. Hence, $f_{t}(I,Q)=\infty$ for any other cases of $I$ and $Q$.
It takes $O(b2^{k})$ time for each introduced check node.
Forget check node: Suppose $t$ is a forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{c\}$, where $c$ is a check node. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(I,Q,d)=\min\{f_{t^{\prime}}(I,Q,d),f_{t^{\prime}}(I\cup\{c\},Q,d-1)\},$ (32)
and
$\begin{split}&g_{t}(I,Q,d)=\mathbb{1}_{f_{t}(I,Q,d)=f_{t^{\prime}}(I,Q,d)}g_{t}(I,Q,d)\\ &+\mathbb{1}_{f_{t}(I,Q,d)=f_{t^{\prime}}(I\cup\{c\},Q,d-1)}g_{t}(I\cup\{c\},Q\{v\},d-1),\end{split}$ (33)
for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$.
Next, we prove that the left-hand side of (32) is greater than or equal to the right-hand side. For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, let $S$ be any trapping set in $G_{t}$ with parameters $(I,Q,d)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S])\cap B_{t^{\prime}}^{c},S\cap B_{t^{\prime}}^{v},\Big{|}\Gamma_{o}(G_{t^{\prime}}[S])\setminus B_{t^{\prime}}^{c}\Big{|}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t}[S])\cap\big{(}B_{t}^{c}\cup\{c\}\big{)},Q,\Big{|}\Gamma_{o}(G_{t}[S])\setminus\big{(}B_{t}^{c}\cup\{c\}\big{)}\Big{|}\Big{)}\\ &=\Big{(}I\cup\big{(}\Gamma_{o}(G_{t}[S])\cap\{c\}\big{)},Q,\\ &\quad\;\Big{|}\big{(}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\big{)}\cap\big{(}\Gamma_{o}(G_{t}[S])\setminus\{c\}\big{)}\Big{|}\Big{)}.\end{split}$
[…]
It takes $O(b2^{k})$ time for each forget check node.
Join node: Suppose $t$ is a join node, and $t_{1}$ and $t_{2}$ be its two child nodes with $B_{t}=B_{t1}=B_{t2}$. Assume we have count $f_{t_{1}}(I,Q,d)$, $g_{t_{1}}(I,Q,d)$, $f_{t_{2}}(I,Q,d)$, and $g_{t_{2}}(I,Q,d)$, for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$. Furthermore, this implies that these functions share the same domain as $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$, which is $(2^{B_{t}^{c}},2^{B_{t}^{v}},\{0,1,...,b\})$. We can compute $f_{t}$ and $g_{t}$ using following equations (34) and (35) for all $I\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$, where (36) takes the minimum over all $(I_{t_{1}},I_{t_{2}},d_{t_{1}},d_{t_{2}})$ satisfying $I=I_{t_{1}}\oplus I_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$ and $d=d_{t_{1}}+d_{t_{2}}$, (37) takes the sum over all $(I_{t_{1}},I_{t_{2}},d_{t_{1}},d_{t_{2}})$ that attain the minimum value in (36) and $\mathbb{1}_{x=y}$ is the indicator function.
$f_{t}(R,Q,d)=\begin{cases}F_{t}(R,Q,d),&\text{if }Q\neq\emptyset;\\ \min\{F_{t}(R,Q,d),f_{t_{1}}(R,Q,d),f_{t_{2}}(R,Q,d)\},&\text{if }Q=\emptyset,\end{cases}$ (34)
$g_{t}(I,Q,d)=\begin{cases}G_{t}(I,Q,d),&\text{if }Q\neq\emptyset;\\ \mathbb{1}_{f_{t}(I,Q,d)=F_{t}(I,Q,d)}G_{t}(I,Q,d)+\mathbb{1}_{f_{t}(I,Q,d)=f_{t_{1}}(I,Q,d)}g_{t_{1}}(I,Q,d)\\ +\mathbb{1}_{f_{t}(I,Q,d)=f_{t_{2}}(I,Q,d)}g_{t_{2}}(I,Q,d),&\text{if }Q=\emptyset,\end{cases}$ (35)
$\begin{split}F_{t}(I,Q,d)=\min_{(I_{t_{1}},I_{t_{2}},d_{t_{1}},d_{t_{2}})}\{f_{t_{1}}(I_{t_{1}},Q,d_{t_{1}})+f_{t_{2}}(I_{t_{2}},Q,d_{t_{2}})-|Q|\},\end{split}$ (36)
$\begin{split}G_{t}(I,Q,d)=\sum_{(I_{t_{1}},I_{t_{2}},d_{t_{1}},d_{t_{2}})}g_{t_{1}}(I_{t_{1}},Q,d_{t_{1}})\times g_{t_{2}}(I_{t_{2}},Q,d_{t_{2}}).\end{split}$ (37)
Next, we prove that the left-hand side of (34) is greater than or equal to the right-hand side. For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, let $S_{t}$ be any trapping set in $G_{t}$ with parameters $(I,Q,d)$. Let $S_{1},S_{2},S_{t_{1}},S_{t_{2}}$ be defined as in the join node part of Theorem 2 when proving that the left-hand side of (16) is greater than or equal to the right-hand side.
When both $S_{t_{1}}$ and $S_{t_{2}}$ are non-empty sets, let $S_{t_{1}}$ have parameters $(I_{t_{1}},Q,d_{t_{1}})$ in $G_{t_{1}}$, and let $S_{t_{2}}$ have parameters $(I_{t_{2}},Q,d_{t_{2}})$ in $G_{t_{2}}$.
According to Lemma 1 and Theorem 2, we have
$\begin{split}&I=\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\cap B_{t}^{c}\\ &=\Big{(}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\Big{)}\cap B_{t}^{c}\\ &=I_{t_{1}}\oplus I_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}]),\end{split}$
and
$\begin{split}&d=\Big{|}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\setminus B_{t}^{c}\Big{|}\\ &=\Big{|}\Big{(}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\Big{)}\setminus B_{t}^{c}\Big{|}\\ &=\Big{|}\Big{(}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\setminus B_{t_{1}}^{c}\Big{)}\oplus\Big{(}\Gamma_{o}(G_{t_{2}}[S_{t_{2}}])\setminus B_{t_{2}}^{c}\Big{)}\\ &\quad\;\oplus\Big{(}\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\setminus B_{t}^{c}\Big{)}\Big{|}\\ &\overset{Prop.2}{=}\Big{|}\Big{(}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\setminus B_{t_{1}}^{c}\Big{)}\cup\Big{(}\Gamma_{o}(G_{t_{2}}[S_{t_{2}}])\setminus B_{t_{2}}^{c}\Big{)}\Big{|}\\ &=d_{t_{1}}+d_{t_{2}}.\end{split}$
This implies that $f_{t}(R,Q,d)\geq F_{t}(R,Q,d)$.
When $S_{t_{1}}=\emptyset$ and $S_{t_{2}}\neq\emptyset$, $S\subset(\cup_{i\in T_{t_{2}}}B_{i})$ is also a trapping set in $G_{t_{2}}$, which implies that the parameters of $S$ in $G_{t_{2}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t_{2}}[S])\cap B_{t_{2}}^{c},S\cap B_{t_{2}}^{v},\Big{|}\Gamma_{o}(G_{t_{2}}[S])\setminus B_{t_{2}}^{c}\Big{|}\Big{)}\\ &\overset{Theorem\;2}{=}\Big{(}\Gamma_{o}(G_{t}[S])\cap B_{t}^{c},Q,\Big{|}\Gamma_{o}(G_{t}[S])\setminus B_{t}^{c}\Big{|}\Big{)}\\ &=(I,Q,d).\end{split}$
Therefore, $f_{t}{(R,Q,d)}\geq f_{t_{2}}(R,Q,d)$.
When $S_{t_{1}}\neq\emptyset$ and $S_{t_{2}}=\emptyset$, a similar proof shows that $f_{t}(R,Q,d)\geq f_{t_{1}}(R,Q,d)$.
By combining these three cases, we can prove that the left-hand side of Equation (34) is greater than or equal to the right-hand side.
For each $(I\subset B_{t}^{c},Q\subset B_{t}^{v},d\in\{0,1,...,b\})$, it takes $O(b2^{k})$ time to obtain the minimum value in (18). Therefore, the computational complexity for each join node is $O(b^{2}2^{2k})$, while for each other type of node, it is $O(b2^{k})$. Given that a nice tree decomposition $(T,(B_{t})_{t\in T})$ of the Tanner graph $G$ contains $O(n)$ nodes in $T$, the overall time complexity of the algorithm is $O(nb^{2}2^{2k})$. ∎
Similar to Section III, we apply the algorithm from Theorem 5 to find an $(a,b)$-trapping set with the minimum $a$ in linear time complexity. Furthermore, we can derive the computational complexity of determining the size and the number of the smallest trapping sets with a given $b>0$ under bounded treewidth constraints.
Corollary 6.
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. For a give $b>0$, we can find an $(a,b)$-trapping sets with minimum $a$ in $G$ in time $O(nb^{2}2^{2k})$.
Corollary 7.
Let $G$ be a Tanner graph with path-width $pw(G)=k^{\prime}$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest trapping sets with $b>0$ in $G$ can be determined in time $O(nb2^{k^{\prime}})$.
V simulation Results
We provide the pseudocode for the algorithm that computes the size and the number of the smallest trapping sets with a given $b$ based on the algorithm in Theorem 5, which includes the case of $b=0$, in the appendix. In this section, we successfully apply the algorithm to a larger number of spatially coupled LDPC codes, for which a tree decomposition, or even a path decomposition, can be easily found. To verify the correctness of the algorithm, we apply it to short spatially coupled LDPC codes, where all small trapping sets can be enumerated using an exhaustive search algorithm.
Figure 4 presents a randomly generated spatially coupled LDPC code, where the uncoupled subcode corresponds to a $3\times 4$ matrix. The coupling length is 10, and the coupling width (as defined in [35]) is 2. Let $G$ denote the Tanner graph corresponding to the spatially coupled LDPC code. We can easily obtain a path decomposition with a width equal to the product of the number of check nodes in the uncoupled subcode and the coupling width plus one. This is achieved by placing each check node along with all its neighboring variable nodes into the same bag, as illustrated in Figure 5.
By applying the algorithm in Theorem 5 to the graph $G$, we obtain that the size of the smallest trapping set with $b=0$ and $b=1$ is 4 and 3, respectively, in $G$. Additionally, the number of trapping sets with $b=0$ and $b=1$ is 7 and 10 in $G$ respectively. To verify the correctness of these results, we perform an exhaustive search algorithm on $G$ to identify all trapping sets with $b=0$ and $a<4$, as well as all trapping sets with $b=1$ and $a<3$. The results are listed in Table I, demonstrating the accuracy of our proposed algorithm.
TABLE I: All the $(a,b)$-trapping set with $a<4,b=0$ and $a<3,b=1$
number (4,0)(4,0)TS in $G$ (3,1)(3,1)TS in $G$
1 13,14,16,20 1,6,7
2 13,16,18,22 2,3,6
3 14,18,20,22 2,4,6
4 25,27,28,30 2,4,7
5 29,30,31,32 2,6,12
6 29,31,34,35 4,5,9
7 30,32,34,35 4,5,12
8 5,7,12
9 17,26,28
10 21,23,26
Figure 4: A randomly generated spatially coupled LDPC code, where the uncoupled subcode corresponds to a $3\times 4$ matrix. The coupling length is 10, and the coupling width is 2. The variable nodes have a regular degree of 3. The black regions indicate positions in the parity-check matrix where the entries are 1, while other positions contain 0.
Figure 5: A path decomposition of the Tanner graph corresponding to Fig.4 with a width of 10. This indicates that the treewidth and pathwidth of graph $G$ are at most 10.
VI Conclusion
For a given $b$, we present an algorithm to compute the size and the number of the smallest trapping sets with $b\geq 0$. For a Tanner graph $G$ with treewidth $k$, the algorithm runs in $O(b^{2}2^{2k})$ time, while for a Tanner graph $G$ with pathwidth $k$, it runs in $O(b2^{k})$ time. We have validated the correctness of the algorithm both theoretically and through simulation experiments. Furthermore, this algorithm can also be used to find an $(a,b)$-trapping set with the minimum $a$.
Appendix A Algorithm
We provide the pseudocode for the algorithm that computes the size and the number of the smallest trapping sets with a given $b$ based on the algorithm in Theorem 5, which includes the case of $b=0$.
Given the Tanner graph $G$ and a rooted nice tree decomposition $(T,(B_{t})_{t\in T})$ of $G$, for any node $t\in T$ and a given $b$, the algorithm does not compute all $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$ over the entire domain $(2^{B_{t}^{c}},2^{B_{t}^{v}},\{0,1,\dots,b\})$. Instead, the values of $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$ are only computed when a trapping set with parameters $(I,Q,d)$ is found in $G_{t}$. In the actual implementation, an array or a linked list $O_{t}$ is used to store those tuples $(I,Q,d)$ for which a corresponding trapping set has been identified. The functions $f_{t}$ and $g_{t}$ can be viewed as arrays or linked lists of the same size as $O_{t}$, indexed by the elements stored in $O_{t}$.
As Algorithm 1, accessing the nodes in the tree decomposition following the order of $\mathcal{T}$ ensures that when computing the functions $f_{t}$ and $g_{t}$ for the current node, the corresponding functions $f_{t^{\prime}}$ and $g_{t^{\prime}}$ for its child nodes have already been computed.
Data: Tanner Graph $G$ and it’s a rooted nice tree decomposition $(T,(B_{t})_{t\in T})$ with root $r$, constant $b$
Result: $f_{r}(\emptyset,\emptyset,b),g_{r}(\emptyset,\emptyset,b)$
1 Sort the nodes in $T$ and denote the sorted sequence as $\mathcal{T}$, such that for any node $t$, its children appear before $t$ in the sequences;
2 foreach $t\in\mathcal{T}$ do
3 if $t$ has a child node then
4     Let $t^{\prime}$ be the child of $t$, then $O_{t^{\prime}}$, $f_{t^{\prime}}$, and $g_{t^{\prime}}$ have already been computed;
5   end if
6 if $t$ has two child nodes then
7     Let $t_{1},t_{2}$ be the child of $t$, then $O_{t_{1}},$ $f_{t_{1}}$, $g_{t_{1}}$, $O_{t_{2}}$, $f_{t_{2}}$, and $g_{t_{2}}$ have already been computed;
8   end if
9
10 switch the type of $t$ do
11    case leaf node do
12       $O_{t}=\emptyset,f_{t}=\emptyset,g_{t}=\emptyset$;
13
14    case introduced variable node do
[…]
16         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
17
18    case forget variable node do
[…]
20         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
21
22    case introduced check node do
[…]
24         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
25
26    case forget check node do
[…]
28         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
29
30    case join node do
[…]
32         delete $O_{t_{1}},f_{t_{1}},g_{t_{1}},O_{t_{2}},f_{t_{2}}$ and $g_{t_{2}}$;
33
34      end case
35
36   end switch
37
38 end foreach
Algorithm 1 computes the size and the number of the smallest trapping sets with a given $b\geq 0$ in $G$.
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1
2Let $v=B_{t}\setminus B_{t^{\prime}}$;
3 $O_{t}=O_{t^{\prime}}$; $f_{t}=f_{t^{\prime}};g_{t}=g_{t^{\prime}}$;
4 $I_{v}=\Gamma_{o}(G[v\cup B_{t}^{c}])$ ;
5 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}\setminus\{(\emptyset,\emptyset,0)\}$ do
6 $(I,Q,d)=(I_{t^{\prime}}\oplus I_{v}),Q_{t^{\prime}}\cup\{v\},d_{t^{\prime}})$;
7 $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
8 $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})+1$;
9 $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
10
11 end foreach
[…]
13 $f_{t}(I_{v},\{v\},0)=1,g_{t}(I_{v},\{v\},0)=1$ ;
Algorithm 2 IVN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $v=B_{t^{\prime}}\setminus B_{t}$;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4
5 if $v\notin Q_{t^{\prime}}$ then
6    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}}\setminus\{v\},d_{t^{\prime}})$
9   end if
10 if $(I,Q,d)\notin O_{t}$ then
11    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
12    $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
13    $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14
15 else
16    $r=\min\{f_{t}(I,Q,d),f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\}$ ;
17    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})}g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
18    $f_{t}(I,Q,d)=r\}$;
19
20   end if
21
22 end foreach
Algorithm 3 FVN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $c=B_{t}\setminus B_{t^{\prime}}$ ;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4 if $c$ has an odd number of edges connected to $Q_{t^{\prime}}$ in $G_{t}$ then
5    $(I,Q,d)=(I_{t^{\prime}}\cup\{c\},Q_{t^{\prime}},d_{t^{\prime}})$;
6
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
9
10   end if
11 $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
12 $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
13 $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14
15 end foreach
Algorithm 4 ICN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $c=B_{t^{\prime}}\setminus B_{t}$ ;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4 if $c\in I_{t^{\prime}}$ then
5    $(I,Q,d)=(I_{t^{\prime}}\setminus\{c\},Q_{t^{\prime}},d_{t^{\prime}}+1)$;
6
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
9
10   end if
11 if $d\leq b$ then
12    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
13    $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14    $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
15
16   end if
17
18 end foreach
Algorithm 5 FCN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t_{1}},f_{t_{1}},g_{t_{1}},O_{t_{2}},f_{t_{2}},g_{t_{2}}$
Result: $O_{t},f_{t},g_{t}$
1 foreach $(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\in O_{t_{1}}$ do
2 foreach $(I_{t_{2}},Q_{t_{2}},d_{t_{2}})\in O_{t_{2}}$ such that $Q_{t_{1}}=Q_{t_{2}}$ do
3    $(I,Q,d)=(I_{t_{1}}\oplus I_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q_{t_{1}}\cup B_{t}^{c}]),Q_{t_{1}},d_{t_{1}}+d_{t_{2}})$;
4    $f^{\prime}(I,Q,d)=f_{t_{1}}(I_{t_{1}},Q_{t_{1}},d_{t_{1}})+f_{t_{2}}(I_{t_{2}},Q_{t_{2}},d_{t_{2}})-|Q|$;
5    $g^{\prime}(I,Q,d)=g_{t_{1}}(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\times g_{t_{2}}(I_{t_{2}},Q_{t_{2}},d_{t_{2}})$;
6    if $(I,Q,d)\notin O_{t}$ then
7       $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
8       $f_{t}(I,Q,d)=f^{\prime}(I,Q,d)$;
9       $g_{t}(I,Q,d)=g^{\prime}(I_{,}Q,d)$;
10
11    else
12
13       $r=\min\{f_{t}(I,Q,d),f^{\prime}(I,Q,d)\}$;
14       $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f^{\prime}(I,Q,d)}g^{\prime}(I,Q,d)$;
15       $f_{t}(I,Q,d)=r$;
16
17      end if
18
19   end foreach
20
21 end foreach
22foreach $(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\in O_{t_{1}}$ such that $Q_{t_{1}}=\emptyset$  do
23 $(I,Q,d)=(I_{t_{1}},\emptyset,d_{t_{1}})$;
24 if $(I,Q,d)\notin O_{t}$ then
25    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
26    $f_{t}(I,Q,d)=f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
27    $g_{t}(I,Q,d)=g_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
28
29 else
30
31    $r=\min\{f_{t}(I,Q,d),f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})\}$;
32    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})}g_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
33    $f_{t}(I,Q,d)=r$;
34
35   end if
36
37 end foreach
38
39foreach $(I_{t_{2}},Q_{t_{2}},d_{t_{2}})\in O_{t_{2}}$ such that $Q_{t_{2}}=\emptyset$ do
40 $(I,Q,d)=(I_{t_{2}},\emptyset,d_{t_{2}})$;
41 if $(I,Q,d)\notin O_{t}$ then
42    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
43    $f_{t}(I,Q,d)=f_{t_{2}}(I_{t_{1}},\emptyset,d_{t_{2}})$;
44    $g_{t}(I,Q,d)=g_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})$;
45
46 else
47
48    $r=\min\{f_{t}(I,Q,d),f_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})\}$;
49    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})}g_{t_{2}}(I_{t_{1}},\emptyset,d_{t_{2}})$;
50    $f_{t}(I,Q,d)=r$;
51
52   end if
53
54 end foreach
[…]